OpenStudy (anonymous):

What is the least common denominator for the fractions 1.6 and 3.4 ? a. 8 B. 12 C. 18 D. 24

OpenStudy (anonymous):

I think its 24

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

The fractions \[\large \sf \frac{1}{6}~and~\frac{3}{4}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well you are somewhat right. 24 is a common denominator, but it isn't the lowest one.

OpenStudy (anonymous):

12? im sorry if it takes me forever to get the answer, math isint my best subject..

OpenStudy (mathmale):

I see you've multiplied 6 and 4 to get a common denominator. This always works, BUT the result is not necessarily the LEAST common denominator. Another way to arrive at the LCD would be to FACTOR each denominator. 4 factors into 2*2 and 6 factors into 3*2. Question: Which factor is common to both denominators? The LCD is found by taking that factor and multiplying it by each of the remaining factors (which here are 2 and 3). LCD=2*2*3=12.

OpenStudy (anonymous):

so was I correct????????? with 12

OpenStudy (mathmale):

Yes, but the question is HOW will you do this next time? Just by multiplying the denomiators or by trying up front to find the LEAST common den.?

OpenStudy (anonymous):

by trying up front to find the least common den. and thanks!